Consider the following hypotheses tests involving the χ2-distribution.

(a) Determine the p-value for Ho: P(1) = P(2) = P(3) = P(4) = 0.25, with χ2 = 9.08. (Give your answer bounds exactly.)
< p <

(b) Determine the p-value for Ho: P(I) = 0.25, P(II) = 0.40, P(III) = 0.35, with χ2 = 9.42. (Give your answer bounds exactly.)
Answer #1

Solution :

(a)

df = 4 - 1 = 3

P-value = P(3 > 9.08) = 0.0282

0.025 < p < 0.05

(b)

df = 3 - 1 = 2

P-value = P(2 > 9.42) = 0.009

P-value < 0.01